please help me answer this equation using slope intercept form x-y=1

-x

Found 2 solutions by larry0304, Alan3354:
Answer by larry0304(1)


Standard slope intercept format is : y = m x + b
Convert the given equation equation to this format :

x-y =1 -------------- : x-y = 1
-x = -x
---------------------
-y = 1 -x
Multiply both side by -1 : -1(-y = 1 -x )
y = -1 + x
Rewrite the equation : y = x -1
Substitute when y=0 , x =? : 0 = x -1
+1 = +1
------------
x = 1
Substitute when y = 1 , x=? : 1 = x -1
Add +1 to both side to : +1 = +1
--------------
x = 2
Replace these values in original
equation if they hold true : x-y =1
when y=0 , x= 1 , check : 1-0 =1 YES
x-y =1
when y=1 , x= 2 , check : 2-1 =1 YES

Answer by Alan3354(69443)

please help me answer this equation using slope intercept form x-y=1

-x
------------
What's the -x all by itself?
----
We don't "answer equations", we solve them for something.
You have one eqn and 2 variables --> an infinite # of "solutions", no unique solution.
-----------
If you graph it, you get a straight line, and every point on it is a solution.